OpenStudy (therealmeeeee):

Please help me I will fan you and give you e medal Solve by the linear combination method (with or without multiplication). 4x – 5y = –18 3x + 2y = –2 A. (–2, 2) B. (–1, 3) C. (–2, 7) D. (3, –1)

OpenStudy (anonymous):

ok what do you think it is?

OpenStudy (anonymous):

i cant help you if you dont want to try

OpenStudy (therealmeeeee):

C

OpenStudy (therealmeeeee):

@jigglypuff314

OpenStudy (anonymous):

no its not c

OpenStudy (therealmeeeee):

B

OpenStudy (anonymous):

each answer is x and y so use those and find the answer to 18 3x + 2y

OpenStudy (therealmeeeee):

ohhhhhh a

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

no i believe its d

OpenStudy (therealmeeeee):

no its a

OpenStudy (anonymous):

what did you get for a

OpenStudy (therealmeeeee):

4g + 5h = 4 8g – 10h = 4 Answer 3: (3/4, 1/5)

OpenStudy (therealmeeeee):

Solve by the linear combination method (with or without multiplication). 4g + 5h = 4 8g – 10h = 4

OpenStudy (anonymous):

ok then yes i wrote one wrong thing and messed it up so yeah that is rigth

OpenStudy (therealmeeeee):

Solve by the linear combination method (with or without multiplication). 4g + 5h = 4 8g – 10h = 4